OpenStudy (anonymous):

4tan^4x+tan^2x-3=sec^2x(4tan^2x-3) verify

OpenStudy (anonymous):

let x= tan^2x and then solve

OpenStudy (anonymous):

you cant use both sides of equation

Directrix (directrix):

4tan^4x+tan^2x-3=sec^2x(4tan^2x-3) ----------------------------------- Factor the left side. 4*tan^4x+tan^2x-3 = (4 tan² x - 3 ) * (tan²x + 1) ========== On the right side, sec^2x(4tan^2x-3), substitute (1+ tan² x) for sec²x to get: (1+ tan² x) * (4*tan² x-3) = (4*tan² x-3) * (tan² x + 1) Now, compare the left and right sides.